OpenStudy (horsegal244):

Simplify 8/-4 ÷ -3/9. Ill give medals

OpenStudy (anonymous):

the answer is 2

OpenStudy (horsegal244):

no negative?

OpenStudy (butterflydreamer):

Firstly, simplify both of your fractions :) 8/-4 = ? -3/9 = ?

OpenStudy (anonymous):

Fraction: 8 / 4 = 23 / 22 = (23 : 22) / (22 : 22) = (8 : 4) / (4 : 4) = 2 / 1 = 2

OpenStudy (horsegal244):

ok ty

OpenStudy (horsegal244):

that is not the answer

OpenStudy (anonymous):

\[\frac{ 8 }{ -4 } ÷ \frac{ -3 }{ 9 }\]

OpenStudy (anonymous):

two negative are = two a psotive so the answer is 2

OpenStudy (horsegal244):

no its not

OpenStudy (anonymous):

its not 2 because 8/-4 = -2 and −3/9 = -1/3

OpenStudy (horsegal244):

these are the answers given

OpenStudy (butterflydreamer):

... 8/-4 = -2 -3/9 = 1/3 You are solving: